OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. 6|6-4x|=8x+4

OpenStudy (oaktree):

Are you there, @mrivera3113?

OpenStudy (anonymous):

Yes I have no idea how to this question I have been stuck on it for 3 days now. @OakTree

OpenStudy (oaktree):

Okay. So we want to start by dividing by 6. What would you get?

OpenStudy (oaktree):

@mrivera3113? Do you know?

OpenStudy (anonymous):

what do I divide 6 to?

OpenStudy (oaktree):

Both sides of the equation.

OpenStudy (anonymous):

I got |6-4x|=8x+.6666

OpenStudy (oaktree):

Not quite. It should be\[\left| 6-4x \right| = \frac{ 4 }{ 3 }x + \frac{ 2 }{ 3 }\]Do you see why?

OpenStudy (oaktree):

Should I give you the answers? Or do you want the explanation?

OpenStudy (anonymous):

Explanation

OpenStudy (oaktree):

Okay. So now that you have the equation above you would write\[6-4x = \frac{ 4 }{ 3 }x + \frac{2}{3}, 6-4x = -\frac{4}{3}x - \frac{2}{3}\]And solve both of those. Do you see why?

OpenStudy (anonymous):

Yes

OpenStudy (oaktree):

Great. So can you solve now?

OpenStudy (anonymous):

Working it out right now.

OpenStudy (oaktree):

Great. Get back to me when you have the answers - I'll tell you if you're right.

OpenStudy (anonymous):

Can I have the explanation on how to do it and answer please.

OpenStudy (oaktree):

The answers are \[x = {1, \frac{ 5 }{ 2 }}\]And as for the how, it's solving algebraic equations. You know, inverse operations?

OpenStudy (anonymous):

Yes i do.

OpenStudy (oaktree):

Great. So you get it?

OpenStudy (anonymous):

Yes thank you soo much!

OpenStudy (oaktree):

No problem. Glad to help.